1.Two columns of the same dataset measuring similar things with the same measurement units can occasionally be tested using a __________ sample t-test.  

2.

Below is an example of what?

randomly selected household objects were found and the prices were placed into a dta matrix

Walmart Amazon

1 $57.48 $54.99

2 $21.19 $24.22

3 $36.36 $36.18

4 $5.90 $7.15

3.What T value would I use to make a 99% confidence interval if sigma is unknown and I have a sample size 25?

4.What T value would I use to make a 90% confidence interval if sigma is unknown and I have a sample size 47?
Answer #1

1.Two columns of the same dataset measuring similar things with the same measurement units can occasionally be tested using a Paired sample t-test

2) This is the example of two sample t test

3) sigma is unknown hence we use t test

n= 25 and alpha = 0.01

Using t table

Tcrit= t(n-1)=t(24,0.005)= 2.7969

4) sigma is unknown hence we use t test

n= 47 alpha= 0.1

From t table

Tcrit= t(n-1)= t(46,0.05)= 1.6787